# UniGaussian: Driving Scene Reconstruction from Multiple Camera Models via Unified Gaussian Representations

### v2

@@ 4.1 Fisheye Rendering Geometric Error Analysis
We employ the method depicted in Fig. 4 to analyze the rendering geometric error introduced by these approximations. First, a 3D Gaussian model is trained using the images of a pinhole camera. Then, images of a fisheye camera are rendered in two different ways: the “convert-project” method as introduced above and the “project-convert” method that renders a pinhole image and distorts it to a fisheye image. The “project-convert” method does not have any approximation in the conversion part, so its rendering result is considered as the reference image. We use the “BICYCLE” and “GARDEN” of the Mip-NeRF360 dataset [1] in this experiment because these data have enough multi-view images and accurate camera poses that enable rendering from any point of view.
@@
The results are presented in Tab. 1. From these results, we can see that the stretching of 3D Gaussians is effective for improving the quality of rendering. This proves the correctness of our rendering algorithm’s derivation because the approximations do not degrade the image quality. Moreover, in our method, most of the rendering time is spent on affine transformations of 3D Gaussians, while rasterization only takes around 3 ms. Note that, the time consumption shown in Tab. 1 is for images with large resolutions, but for autonomous driving scenes, 3D Gaussians are projected within a frustum and the image resolution is less than $4358\times 2824$ so the rendering can be done in real time. Besides, although the 2nd order approximation of the polar stretching ratio gives an improvement on rendering quality, it greatly increases the rendering time.
@@ 4.2 Driving Scene Fisheye Camera Simulation
In this experiment, we verify the effectiveness of the proposed fisheye rendering method on the KITTI-360 dataset [15]. We compare the image quality of both global image and three representative local zones, namely, A, B, and C, as illustrated in Fig. 5. Details of the experimental setup can be found in the supplementary material.
We present the quantitative results in Tab. 2 and the qualitative results in Fig. 6. Overall, our approach achieves the best rendering quality both globally and locally (at zones A, B and C), while maintaining real-time rendering speed. In comparison, 3DGS+Undistort performs poorly on fisheye simulation, which indicates that direct distortion is not a suitable choice. Besides, from Tab. 2 and Fig. 7, we can see that our method without stretching yields slightly worse image quality, especially for texture details, but achieves significantly faster rendering speed of 138 FPS. Also, skipping stretching increases the number of Gaussians by $40\%$, which increases the risk of GPU memory overflow. The reason is that small Gaussians are insensitive to changes in the distribution caused by light rays distortion and the optimizer tends to approximate the scene with a large number of small 3D Gaussians in exchange for better rendering quality.
@@
Instant-NGP[21] 24.6 0.808 0.181 - - -
UniSim-SF[32, 25] 24.9 0.812 0.184 - - -
AlignMiF[25] 25.3 0.826 0.164 - - -
@@
Table 3: Results of multiple camera model simulation on KITTI-360. “HUGS++” denotes using our method for additionally training fisheye cameras. Some results are borrowed from [25].
@@ 4.3 Multiple Camera Model Simulation
In this experiment, we evaluate our UniGaussian framework for driving scene simulation from multiple camera models. We conduct our experiments on KITTI-360 [15], because it is a real-world autonomous driving dataset that provides both pinhole and fisheye images, while the other commonly used datasets provide only one type of image.
As shown in Tab. 3, for pinhole camera simulation, for fair comparison with AlignMiF and UniSim-SF, we also report the results of our approach with pinhole camera only. We can see that both ours and ours(pinhole) achieve better results than the state-of-the-art driving scene reconstruction methods, such as HUGS, AlignMiF and UniSim-SF. As for fisheye camera simulation, we add our fisheye rendering method to HUGS (named HUGS++) because HUGS does not support fisheye cameras. The results in Tab. 3 show that our approach outperforms the modified method. Besides, as shown in Fig. 8, our approach generates better rendering images with more fine-grained details. For example, our approach renders the license plate, the fence, the taillight, etc., while HUGS++ generates worse rendering images. Moreover, we also visualize some semantic and depth maps in Fig. 8. We can see that our multimodality outputs achieve slightly better holistic 3D scene understanding.
@@ 5 Conclusion
This supplementary material is organized as follows: In Sec. 6, we present more details of the experimental setups; in Sec. 7, we provide the implementation details of our approach; in Sec. 8, we present the experimental comparison with Fisheye-GS [16]; in Sec. 9, we present results of novel view synthesis with lane shift; in Sec. 10, we discuss the details of our approach to Lidar simulation; in Sec. 11, we discuss the limitation and future work of our approach.
@@ 6.1 Experimental Setup for Fisheye Rendering Geometric Error Analysis
The pinhole and fisheye cameras are set to have the same FOV and resolution during rendering. To ensure the rendering quality of the “project-convert” method, the FOV is set close to the FOV of the training data. The resolution needs to be set to a large value because the pinhole image will shrink to the center area after distortion. The 3DGS hyper-parameters are mainly set following [10].
@@ 6.2 Experimental Setup for Driving Scene Fisheye Camera Simulation
We conduct experiments on the KITTI-360 dataset [15]. KITTI-360 contains images captured from fisheye cameras in real-world driving scenes and provides camera calibration and ego-vehicle pose. We select a segment of 221 images (frame ids 227-447 from “2013_05_28_drive_0000_sync”) and evenly select every 8th image as the test set while the others are used as the training set. In addition to comparing the global image quality, three local zones (A, B, C) are compared. As shown in Fig. 5, zone “A” (Blue) is the bottom area where objects appear usually very close to the camera, zone “B” (Green) is the sky with little texture where the geometry is not very accurate, and zone “C” (Red) is far away from the camera with the large distortion in this region.
@@
The MEI camera model is the only model provided for the fisheye camera on KITTI-360, so to evaluate our method on both the MEI and Kannala-Brandt models, we convert the MEI model to the Kannala-Brandt model. Specifically, the conversion is an approximation of $\theta_{d}(\theta)$ determined by Eq. (3) of the main paper with the series expansion Eq. (2) of the main paper. Note that, when $\theta=0$, the first derivative of Eq. (2) equals $1$, while the first derivative of Eq. (3) is $1/(1+\xi)$. To eliminate this difference, Eq. (3) is modified as:
@@
We compare our approach with three NeRF-based methods, including Instant-NGP [21], Nerfacto-big [24], Zip-NeRF [2], and one 3DGS-based method, namely “3DGS [10]+Undistort”. To adapt 3DGS for driving scene reconstruction with fisheye images, we rectangularize fisheye images for reconstruction and render scene with the pinhole camera model and distort to fisheye images (named 3DGS+Undistort). Besides, the NeRF-based methods are implemented based on [24].
@@ 6.3 Experimental Setup for Multiple Camera Model Simulation
We conduct our experiments on KITTI-360 [15], because it is a real-world autonomous driving dataset that provides both pinhole and fisheye images, while the other commonly used datasets provide only one type of image. For evaluation, we select four sequences as [25]. Each sequence contains 64 frames, and we select every 4th frame as the test set while the others are used as the training set. Note that, although there are fisheye datasets for autonomous driving, such as Woodscape [33], they are constructed for perception tasks (e.g., object detection, segmentation, etc.) rather than driving scene reconstruction, so they often only have discrete frames and lack continuous video sequence data. In contrast, KITTI-360 provides a complete real-world autonomous driving sensor suite, including fisheye, pinhole and LiDAR, so evaluation on KITTI-360 for multiple camera model simulation is comprehensive to verify the effectiveness of our approach.
@@
To the best of our knowledge, no existing driving scene reconstruction method simulates both pinhole and fisheye cameras in a unified framework. HUGS [35] is a state-of-the-art 3DGS-based method for driving scene reconstruction and 3D scene understanding. We therefore use it as a baseline method for comparison. We modify HUGS based on the official code and the paper and additionally use our differentiable rendering method to train HUGS for fisheye camera simulation. For the other compared driving scene simulation methods, we borrow the results from [25]. Note that, although there are some other driving scene simulation methods, they are mostly designed for pinhole camera simulation, so we choose the representative HUGS [35], AlignMiF[25] and UniSim-SF[32, 25] for comparison in our experiments. Modifying existing state-of-the-art methods for fisheye camera simulation in autonomous driving is beyond the scope of this work.
@@ 7 Implementation Details
We implement our approach using python and pytorch. Following [10, 35], we set the initial position learning rate to $1.6\times 10^{-4}$ with decay to $1.6\times 10^{-6}$ at the last iteration. Both the scaling and rotation learning rates are set to $0.001$, while the learning rate of the spherical harmonics feature is set to $0.0025$ and the opacity learning rate is set to $0.05$. Adaptive Density Control is performed every 100 iterations starting from 500 iterations. For the composite scene Gaussians, the background Gaussians are initialized with LiDAR point clouds, the dynamic Gaussians are randomly initialized, and the sky Gaussians are uniformly initialized in a distant (more than 100 meters) spherical region for simplicity, but more advanced sky cubemaps [31] can be used. The training loss $\mathcal{L}$ of our approach is defined in Eq. (20) of the main paper. The details of these losses and regularization terms are listed below.
@@
where $\mathcal{L}_{1}$ is the L1 loss, $\mathcal{L}_{SSIM}$ is a D-SSIM term [10], and $\lambda_{rgb}$ is set to 0.2 following [10].
@@
where $\mathcal{L}_{d-lidar}$ is the L1 loss between the depth derived from LiDAR point clouds and the masked region from the rendered depth map, and $\mathcal{L}_{d-mono}$ is the Pearson depth loss [29] between the rendered depth map and the monocular depth map computed with [8].
@@
It is the semantic loss for the rendering semantic map $M_{s}$ and the predefined 2D semantic segmentation map [14]. This is implemented with the cross-entropy loss to classify the semantic logits and its weight is set to 0.01. This loss helps to generate semantic maps for driving scenes, improving the holistic driving scene understanding.
@@
Following [11], we employ the Gaussian opacity and scale regularization term to encourage a compact scene Gaussian representation. It is defined as:
@@ 8 Experimental Comparison with Fisheye-GS
As discussed in Sec. 2.1, Fisheye-GS [16] is only based on ideal camera models with equidistant projection, which hinders its use in driving scene reconstruction because fisheye cameras in driving scene are usually generic models and have large FOVs. Since Fisheye-GS is a state-of-the-art method adapting 3DGS to fisheye cameras, it would be interesting to compare our differentiable fisheye rendering method with Fisheye-GS. To this end, we conduct experimental comparison on Zip-NeRF(fisheye) [2] with four large scenes, namely, Alameda, Berlin, London, and Nyc. These scenes are respectively captured with a fisheye lens of 180 degree. We select every 8th frame as the test set while the others are used as the training set. As shown in Tab. 5, we approach achieves better performance compared with Fisheye-GS. Specifically, in the Alameda, London and Nyc scenes, we approach achieves significantly better results, while in the Berlin scene, our approach achieves comparable performance compared with Fisheye-GS. On average, our approach yields PSNR of 25.0 dB while Fisheye-GS obtains PSNR of 23.4 dB. These results verify the superiority of our approach over Fisheye-GS.
@@ 10 Details of the Optional LiDAR Simulation
where $x_{sim}$ and $I_{sim}$ are the simulated point cloud positions and intensities, $x_{gt}$ and $I_{gt}$ are ground truths, and $\lambda_{l}$ is set to $0.1$. In Fig. 10, we visualize some simulated LiDAR point cloud maps and intensity maps. From these results, we can observe that the simulated point cloud and intensity maps are close to the real LiDAR, especially for the nearby regions. On the other hand, we can also observe some noisy points because the indirect simulation strategy relies heavily on the accuracy of the depth maps. Note that, the LiDAR simulation is not the focus of this work and more effort should be made in order to improve LiDAR simulation with 3DGS and to simulate more complex LiDAR phenomena.

### v3

@@ 4.1 Fisheye Rendering Geometric Error Analysis
We employ the method depicted in Fig. 4 to analyze the rendering geometric error introduced by these approximations. First, a 3D Gaussian model is trained using the images of a pinhole camera. Then, images of a fisheye camera are rendered in two different ways: the “convert-project” method as introduced above and the “project-convert” method that renders a pinhole image and distorts it to a fisheye image. The “project-convert” method does not have any approximation in the conversion part, so its rendering result is considered as the reference image. We use the “BICYCLE” and “GARDEN” of the Mip-NeRF360 dataset [barron2022mip] in this experiment because these data have enough multi-view images and accurate camera poses that enable rendering from any point of view.
@@
The results are presented in Tab. 1. From these results, we can see that the stretching of 3D Gaussians is effective for improving the quality of rendering. This proves the correctness of our rendering algorithm’s derivation because the approximations do not degrade the image quality. Moreover, in our method, most of the rendering time is spent on affine transformations of 3D Gaussians, while rasterization only takes around 3 ms. Note that the time consumption shown in Tab. 1 is for images with large resolutions, but for autonomous driving scenes, 3D Gaussians are projected within a frustum and the image resolution is less than $4358\times 2824$ so the rendering can be done in real time. Besides, although the 2nd order approximation of the polar stretching ratio gives an improvement on rendering quality, it greatly increases the rendering time.
@@ 4.2 Driving Scene Fisheye Camera Simulation
In this experiment, we verify the effectiveness of the proposed fisheye rendering method on the KITTI-360 dataset [liao2022kitti]. We compare the image quality of both global image and three representative local zones, namely, A, B, and C, as illustrated in Fig. 5. Details of the experimental setup can be found in the supplementary material.
We present the quantitative results in Tab. 2 and the qualitative results in Fig. 6. Overall, our approach achieves the best rendering quality both globally and locally (at zones A, B and C), while maintaining real-time rendering speed. In comparison, 3DGS+Undistort performs poorly in fisheye simulation, indicating that direct distortion is not a suitable choice. Besides, from Tab. 2 and Fig. 7, we can see that our method without stretching yields slightly worse image quality, especially for texture details, but achieves significantly faster rendering speed of 138 FPS. Also, skipping stretching increases the number of Gaussians by $40\%$, which increases the risk of GPU memory overflow. The reason is that small Gaussians are insensitive to changes in the distribution caused by light rays distortion and the optimizer tends to approximate the scene with a large number of small 3D Gaussians in exchange for better rendering quality.
@@
Instant-NGP[muller2022instant] 24.6 0.808 0.181 - - -
UniSim-SF[yang2023unisim, tao2024alignmif] 24.9 0.812 0.184 - - -
AlignMiF[tao2024alignmif] 25.3 0.826 0.164 - - -
@@
Table 3: Results of multiple camera model simulation on KITTI-360. “HUGS++” denotes using our method for additionally training fisheye cameras, and for a fair comparison, “HUGS++” uses the same initialization as our approach. Some results are borrowed from [tao2024alignmif].
@@ 4.3 Multiple Camera Model Simulation
In this experiment, we evaluate our UniGaussian framework for driving scene simulation from multiple camera models. We conduct our experiments on KITTI-360 [liao2022kitti], because it is a real-world autonomous driving dataset that provides both pinhole and fisheye images, while the other commonly used datasets provide only one type of image.
As shown in Tab. 3, for pinhole camera simulation, for fair comparison with AlignMiF and UniSim-SF, we also report the results of our approach with pinhole camera only. We can see that both ours and ours(pinhole) achieve better results than the state-of-the-art driving scene reconstruction methods, such as HUGS, AlignMiF and UniSim-SF. As for fisheye camera simulation, we add our fisheye rendering method to HUGS (named HUGS++) because HUGS does not support fisheye cameras. The results in Tab. 3 show that our approach outperforms the modified method. Besides, as shown in Fig. 8, our approach generates better rendering images with more fine-grained details. For example, our approach renders the license plate, the fence, the taillight, etc., while HUGS++ generates worse rendering images. Moreover, we also visualize some semantic and depth maps in Fig. 8. We can see that our multimodal outputs achieve slightly better holistic 3D scene understanding.
@@ 5 Conclusion
This supplementary material is organized as follows: in Sec. 6, we present more details of the experimental setups; in Sec. 7, we provide the implementation details of our approach; in Sec. 8, we present the experimental comparison with Fisheye-GS [liao2024fisheye]; in Sec. 9, we present results of novel view synthesis with lane shift; in Sec. 10, we discuss the details of our approach to Lidar simulation; in Sec. 11, we discuss the key difference between NeRF-based methods and our approach; in Sec. 12, we discuss the limitation and future work of our approach.
@@ 6.1 Experimental Setup for Fisheye Rendering Geometric Error Analysis
The pinhole and fisheye cameras are set to have the same FOV and resolution during rendering. To ensure the rendering quality of the “project-convert” method, the FOV is set close to the FOV of the training data. The resolution needs to be set to a large value because the pinhole image will shrink to the center area after distortion. The 3DGS hyper-parameters are mainly set following [kerbl3Dgaussians].
@@ 6.2 Experimental Setup for Driving Scene Fisheye Camera Simulation
We conduct experiments on the KITTI-360 dataset [liao2022kitti]. KITTI-360 contains images captured from fisheye cameras in real-world driving scenes and provides camera calibration and ego-vehicle pose. We select a segment of 221 images (frame ids 227-447 from “2013_05_28_drive_0000_sync”) and evenly select every eighth image as the test set while the others are used as the training set. In addition to comparing the global image quality, three local zones (A, B, C) are compared. As shown in Fig. 5, zone “A” (Blue) is the bottom area where objects appear usually very close to the camera, zone “B” (Green) is the sky with little texture where the geometry is not very accurate, and zone “C” (Red) is far away from the camera with the large distortion in this region.
@@
The MEI camera model is the only model provided for the fisheye camera on KITTI-360, so to evaluate our method on both the MEI and Kannala-Brandt models, we convert the MEI model to the Kannala-Brandt model. Specifically, the conversion is an approximation of $\theta_{d}(\theta)$ determined by Eq. (3) of the main paper with the series expansion Eq. (2) of the main paper. Note that when $\theta=0$, the first derivative of Eq. (2) equals $1$, while the first derivative of Eq. (3) is $1/(1+\xi)$. To eliminate this difference, Eq. (3) is modified as:
@@
We compare our approach with three NeRF-based methods, including Instant-NGP [muller2022instant], Nerfacto-big [nerfstudio], Zip-NeRF [barron2023zip], and one 3DGS-based method, namely “3DGS [kerbl3Dgaussians]+Undistort”. To adapt 3DGS for driving scene reconstruction with fisheye images, we rectangularize fisheye images for reconstruction and render scene with the pinhole camera model and distort to fisheye images (named 3DGS+Undistort). Besides, the NeRF-based methods are implemented based on [nerfstudio].
@@ 6.3 Experimental Setup for Multiple Camera Model Simulation
We conduct our experiments on KITTI-360 [liao2022kitti], because it is a real-world autonomous driving dataset that provides both pinhole and fisheye images, while the other commonly used datasets provide only one type of image. For evaluation, we select four sequences as [tao2024alignmif]. Each sequence contains 64 frames, and we select every fourth frame as the test set while the others are used as the training set. Note that although there are fisheye datasets for autonomous driving, such as Woodscape [yogamani2019woodscape], they are constructed for perception tasks (e.g., object detection, segmentation, etc.) rather than driving scene reconstruction, so they often only have discrete frames and lack continuous video sequence data. In contrast, KITTI-360 provides a complete real-world autonomous driving sensor suite, including fisheye, pinhole and LiDAR, so evaluation on KITTI-360 for multiple camera model simulation is comprehensive to verify the effectiveness of our approach.
@@
To the best of our knowledge, no existing driving scene reconstruction method simulates both pinhole and fisheye cameras in a unified framework. HUGS [zhou2024hugs] is a state-of-the-art 3DGS-based method for driving scene reconstruction and 3D scene understanding. We therefore use it as a baseline method for comparison. We modify HUGS based on the official code and the paper and additionally use our differentiable rendering method to train HUGS for fisheye camera simulation. For the other compared driving scene simulation methods, we borrow the results from [tao2024alignmif]. Note that although there are some other driving scene simulation methods, they are mostly designed for pinhole camera simulation, so we choose the representative HUGS [zhou2024hugs], AlignMiF[tao2024alignmif] and UniSim-SF[yang2023unisim, tao2024alignmif] for comparison in our experiments. Modifying existing state-of-the-art methods for fisheye camera simulation in autonomous driving is beyond the scope of this work.
@@ 7 Implementation Details
We implement our approach using python and pytorch. Following [kerbl3Dgaussians, zhou2024hugs], we set the initial position learning rate to $1.6\times 10^{-4}$ with a decay to $1.6\times 10^{-6}$ at the last iteration. Both scaling and rotation learning rates are set to $0.001$, the learning rate of the spherical harmonics feature is set to $0.0025$ and the opacity learning rate is set to $0.05$. Adaptive Density Control is performed every 100 iterations starting from 500 iterations. For composite scene Gaussians, the background Gaussians are initialized with LiDAR point clouds, the dynamic Gaussians are randomly initialized, and the sky Gaussians are uniformly initialized in a distant (more than 100 meters) spherical region for simplicity, but more advanced sky cubemaps [yan2024street] can be used. The training loss $\mathcal{L}$ of our approach is defined in Eq. (20) of the main paper. The details of these losses and regularization terms are listed below.
@@
where $\mathcal{L}_{1}$ is the L1 loss, $\mathcal{L}_{SSIM}$ is a D-SSIM term [kerbl3Dgaussians], and $\lambda_{rgb}$ is set to 0.2 following [kerbl3Dgaussians].
@@
where $\mathcal{L}_{d-lidar}$ is the L1 loss between the depth derived from LiDAR point clouds and the masked region from the rendered depth map, and $\mathcal{L}_{d-mono}$ is the Pearson depth loss [xiong2023sparsegs] between the rendered depth map and the monocular depth map computed with [hu2024metric3d].
@@
It is the semantic loss for the rendering semantic map $M_{s}$ and the predefined 2D semantic segmentation map [Liao2022PAMI]. This is implemented with the cross-entropy loss to classify the semantic logits and its weight is set to 0.01. This loss helps to generate semantic maps for driving scenes, improving the holistic driving scene understanding.
@@
Following [kheradmand20243d], we employ the Gaussian opacity and scale regularization term to encourage a compact scene Gaussian representation. It is defined as:
@@ 8 Experimental Comparison with Fisheye-GS
As discussed in Sec. 2, Fisheye-GS [liao2024fisheye] is only based on ideal camera models with equidistant projection, which hinders its use in driving scene reconstruction because fisheye cameras in driving scene are usually generic models and have large FOVs. Since Fisheye-GS is a state-of-the-art method adapting 3DGS to fisheye cameras, it would be interesting to compare our differentiable fisheye rendering method with Fisheye-GS. To this end, we conduct experimental comparison on Zip-NeRF(fisheye) [barron2023zip] with four large scenes, namely, Alameda, Berlin, London, and Nyc. These scenes are respectively captured with a fisheye lens of 180 degree. We select every eighth frame as the test set while the others are used as the training set. As shown in Tab. 5, our approach achieves better performance compared with Fisheye-GS. Specifically, in the Alameda, London and Nyc scenes, our approach achieves significantly better results, while in the Berlin scene, our approach achieves comparable performance compared with Fisheye-GS. On average, our approach yields PSNR of 25.0 dB while Fisheye-GS obtains PSNR of 23.4 dB. These results verify the superiority of our approach over Fisheye-GS.
@@ 10 Details of the Optional LiDAR Simulation
where $x_{sim}$ and $I_{sim}$ are the simulated point cloud positions and intensities, $x_{gt}$ and $I_{gt}$ are ground truths, and $\lambda_{l}$ is set to $0.1$. In Fig. 10, we visualize some simulated LiDAR point cloud maps and intensity maps. From these results, we can observe that the simulated point cloud and intensity maps are close to the real LiDAR, especially for the nearby regions. On the other hand, we can also observe some noisy points because the indirect simulation strategy relies heavily on the accuracy of the depth maps. Note that the LiDAR simulation is not the focus of this work and more effort should be made in order to improve LiDAR simulation with 3DGS and to simulate more complex LiDAR phenomena.
11 Key Difference between NeRF-based methods and our approach
Although there have been some NeRF-based methods for fisheye rendering, such as [choi2024dico], they are mostly time-consuming and cannot be combined with Gaussian splatting for driving scene simulation. Besides, in [choi2024dico], Choi et al. show that the performance of their approach is close to Instant-NGP and NeRFacto, while our experiments in Tab. 2 show that our approach significantly outperforms Instant-NGP and NeRFacto. More importantly, we did not thoroughly compare these NeRF-based methods because the novelty of our approach is a new fisheye method to solve limitation of 3DGS and a unified framework for driving scene reconstruction, while the performance of NeRF-based methods does not diminish the novelty of our approach.